# Aharoni’s rainbow cycle conjecture holds up to an additive constant

### v3

@@ 1. Introduction
In Section 2 we state some preliminary lemmas needed in the proof of our main theorem. Our proof is split into two cases, depending on whether there are many ‘non-star’ vertices or few non-star vertices. This case distinction is described precisely in Section 3, where we also prove a key lemma used in the proof. Then, we handle the many non-star case in Section 4, and the few non-star case in Section 5.
@@ 2. Preliminaries
$g\leqslant\frac{n+r+\mathsf{def}_{r}(D)}{r}+2r^{2}.$
@@
Note that if $n<2r^{2}$, then the statement is immediately true, so we may assume that $n\geq 2r^{2}$. Now, suppose that the statement is false. In particular,
$g>\frac{n+r+\mathsf{def}_{r}(D)}{r}.$
Since $n\geq 2r^{2}$, we have
$g>\frac{n}{r}+1\geqslant 2r-1.$
It follows by Theorem 2.3 that
$g\leqslant\frac{n+\mathsf{def}_{r}(D)-1}{r}+1\leqslant\frac{n+r+\mathsf{def}_{r}(D)}{r},$
which is a contradiction, as desired. This completes the proof. ∎
@@ 3. The Set-up
For the remainder of the paper, $G$ is a simple edge-colored graph with $n$ vertices, $n$ colors, and each color class of size at least $2$ and at most $r$, for $r\geq 2$. In other words, $G$ satisfies the hypotheses of Theorem 1.7.
@@
Let $R=X\cup\{u_{1},\cdots,u_{m}\}\subseteq V(G)$, $M:=\{M_{u_{i}}:i\in[m]\}$ be a galaxy rooted at $\{u_{1},\cdots,u_{m}\}$, $C_{X}\subseteq C(X)$, and $x\in\mathbb{N}$ be such that
@@
    for every pair of vertices $u,v\in X$, there exists a rainbow path consisting of colors in $C_{X}$ from $u$ to $v$ in $G[X]$ of length at most $x$.
@@
$\frac{m+\sum_{i=1}^{m}(r-\lvert u_{i}\rvert_{M})}{r}+x+2.$
@@
See Figure 1 for an example with $m=7$. For each $u\in R$, we define a function $p:R\to R$ as follows. If $u\in X$, then $p(u)=u$. If $u\notin X$, but some $M$-neighbor $v$ of $u$ is in $X$, we arbitrarily choose one such $v$, and define $p(u)=v$. Otherwise, define $p(u)=u_{i}$, where $i$ is the minimum index such that $u_{i}$ is an $M$-neighbor of $u$.
@@
For the remaining case, let $u^{\prime}\in X$ and $v^{\prime}\in X$ be the other ends of $P^{u}$ and $P^{v}$. Let $Q^{u}=P^{u}\setminus u^{\prime}$ and $Q^{v}=P^{v}\setminus v^{\prime}$, and let $S^{u}$ be $Q^{u}$ minus its last vertex and $S^{v}$ be $Q^{v}$ minus its last vertex. By (3), we may assume that at least one of them is not in $X$, namely that $u=u_{a}$ for some $a\in[m]$ and that either $v\in X$ or $v=u_{b}$ for some $b<a$. By the previous case, we may assume that for all $x\in V(Q^{u})$ and $y\in V(Q^{v})$, $x$ and $y$ do not have any common $M$-neighbors in $\{u_{1},\dots,u_{m}\}$. Therefore,
@@
By (3), there is a rainbow path $P$ consisting of colors in $C_{X}$ from $u^{\prime}$ to $v^{\prime}$ in $G[X]$ of length at most $x$. Since $C_{X}\cap C(M)=\emptyset$, $P^{u}\cup P^{v}\cup P$ is a rainbow path consisting of colors in $C_{X}\cup C(M)$ from $u$ to $v$ in $G[R]$ of length at most
$\frac{m+\sum_{i=1}^{m}(r-|u_{i}|_{M})}{r}+x+2,$
@@ 4. Many Non-Stars
Let $k:=\lceil 78r\log{r}\rceil$ and $f(r)=80k^{2}r^{2}$. We will show that, for fixed $r$, we may take $\alpha_{r}:=(4r+2)f(r)+2r^{2}$ in Theorem 1.7. To prove Theorem 1.7, we proceed by induction on $n$. The base case is $n\leq\alpha_{r}$, which is true since there is always a rainbow cycle of length at most $n$.
@@
In this section we consider the case where $|N|>f(r)$. We say that a subset $N^{\prime}\subseteq N$ covers a color class $A$ if every edge in $A$ is incident to at least one vertex in $N^{\prime}$.
@@
subsets of $N$ of size $k$ cover some color class. Therefore, there exists a subset $N^{\prime}\subseteq N$ of size $k$ such that $N^{\prime}$ does not cover any color class. It follows that $G$ has a rainbow subgraph $H$ with $n$ edges and $V(H)\subseteq V(G)\setminus N^{\prime}$. In particular, $H$ is excess-$k$. Since $r\geq 2$ and $n\geq k$ (otherwise we have a rainbow cycle of length at most $k<\alpha_{r}$),  Corollary 2.2 gives that $H$ has a rainbow cycle of length at most
@@
Now, let $u_{1},\cdots,u_{m}$ be a maximal sequence of vertices in $G\setminus X$ such that there is a star color class $A_{i}$ centred at $u_{i}$ with $V(A_{i})\setminus\{u_{i}\}\subseteq X\cup\{u_{1},\dots,u_{i-1}\}$. Let $T^{\prime}=X\cup\{u_{1},\cdots,u_{m}\}$, $c_{i}$ denote the color of $A_{i}$, and $C^{\prime}(T^{\prime})=C(X)\cup\{c_{1},\cdots,c_{m}\}$. By Lemma 3.1, for all $u,v\in T^{\prime}$ there exists a rainbow path in $G[T^{\prime}]$ from $u$ to $v$ consisting of colors in $C(X)\cup\{c_{1},\dots,c_{m}\}$ of length at most
$\frac{m+\sum_{i=1}^{m}(r-\lvert u_{i}\rvert_{M})}{r}+x+2$ $=\frac{m+\sum_{i=1}^{m}(r-\lvert u_{i}\rvert_{M})}{r}+\left(\frac{|X|+\sum_{c\in C(X)}(r-|c_{G}|)}{r}-5r+2\right)+2$
@@
Therefore, $|\mathcal{A}|\leqslant 4r-1$; otherwise, $T^{\prime}\cup Y$ contradicts the maximality of $X$. Now, we contract $T^{\prime}$ to a single vertex, delete all edges with colors in $T^{\prime}$, and remove parallel edges of the same color. Let the resulting edge-colored graph be $G^{\prime}$. Note that by the construction of $T^{\prime}$, $G^{\prime}$ contains $|V(G^{\prime})|$ color classes of size at least $2$, and at most $14r-$ 1 color classes have less edges in $G^{\prime}$ than they do in $G$ (since $|\mathcal{A}|\leqslant 4r-1$). By induction on $n$, $G^{\prime}$ has a rainbow cycle $D$ of length at most
@@ 5. Few Non-Stars
Let $N=\{v_{1},v_{2},\cdots,v_{t}\}$ be an arbitrary fixed ordering of the vertices of $N$. Let $U=V(G)\setminus N$ and choose a galaxy $M:=(M_{u}:u\in U)$ rooted at $U$ by arbitrarily choosing a star class $M_{u}$ of color $c_{u}$ centred at $u$ for each $u\in U$. An edge $e\in E(G)$ is an $M$-edge if $e\in E(M_{u})$ for some $u\in U$. For $Y\subseteq V(G)$ we define $\gamma(Y)$ to be the number of star vertices $u\in V(G)\setminus Y$ such that at least one $M$-neighbor of $u$ is in $Y$.
@@
    $\sum_{i=1}^{j}d_{i}+\gamma(\bigcup_{i=1}^{j}T_{i})\leqslant(4r+1)j$.
@@
By construction, clearly $T_{j}$ satisfies Item 1 and Item 2. Let $C^{\prime}:=\{c_{u_{i}}:i\in[m]\}$. By Lemma 3.1, for any pair of vertices $u,v\in T_{j}$, there is a rainbow path of $M$-edges from $u$ to $v$ in $G[T_{j}]$ of length at most
$\frac{m+\sum_{c\in C^{\prime}}(r-|c_{G}|)}{r}+x+2$ $=\frac{m+\sum_{c\in C^{\prime}}(r-|c_{G}|)}{r}+\left(\frac{|X|+\sum_{c\in C_{M}(X)}(r-|c_{G}|)}{r}+|s(X)|\right)+2$
@@
$\sum_{i=1}^{j}d_{i}+\gamma(\bigcup_{i=1}^{j}T_{i})\leqslant(4r+1)j$.
@@
If $|Y|\geqslant 4r$, then $Y\cup T_{j}$ contradicts the maximality of $X$. Thus, $|Y|\leqslant 4r-1$. Observe that $\gamma(\bigcup_{i=1}^{j}T_{i})\leqslant|Y|-|s(T_{j})|+\gamma(\bigcup_{i=1}^{j-1}T_{i})$. Therefore,
$\sum_{i=1}^{j}d_{i}+\gamma(\bigcup_{i=1}^{j}T_{i})$ $\leqslant d_{j}+|Y|-|s(T_{j})|+\sum_{i=1}^{j-1}d_{i}+\gamma(\bigcup_{i=1}^{j-1}T_{i})$
$=2+|Y|+\sum_{i=1}^{j-1}d_{i}+\gamma(\bigcup_{i=1}^{j-1}T_{i})$
$\leqslant(4r+1)+\sum_{i=1}^{j-1}d_{i}+\gamma(\bigcup_{i=1}^{j-1}T_{i})$
@@
The remaining case is $V(G)\neq\bigcup_{i=1}^{t}T_{i}$. Let $W=G\setminus\left(\bigcup_{i=1}^{t}T_{i}\right)$. By Item 2, every vertex $v\in W$ has at least one of its $M$-neighbors in $W$. Note that by Item 4, at most $(4r+1)f(r)$ vertices in $W$ have an $M$-neighbor outside of $W$.
Now, let $D$ be a digraph with vertex set $W$ and for each $v\in W$ and each $M$-neighbor $u$ of $v$ in $W$, put an arc from $v$ to $u$ in $D$. Then we apply Corollary 2.4 in $D$ to obtain a directed cycle $C$ in $D$ of length at most
$\frac{|W|+r+\sum_{c\in C(G)}(r-|c_{G}|)+(4r+1)(r-1)f(r)}{r}+2r^{2}$ $\leqslant\frac{n+\sum_{c\in C(G)}(r-|c_{G}|)}{r}+\alpha_{r}.$

### v4

@@ 1. Introduction
In Section 2 we state some preliminary lemmas needed in the proof of our main theorem. Our proof is split into two cases, depending on whether there are many ‘non-star’ vertices or few non-star vertices. We set-up our notation and assumptions in Section 3, where we also prove a key lemma used in the proof. Then, we handle the many non-star case in Section 4, and the few non-star case in Section 5.
@@ 2. Preliminaries
$g\leqslant\frac{n+\mathsf{def}_{r}(D)}{r}+2r-1.$
@@
The statement is obviously true if $g\leqslant 2r-2$. Thus, we may assume that $g\geqslant 2r-1$. It follows by Theorem 2.3 that
$g\leqslant\frac{n+\mathsf{def}_{r}(D)-1}{r}+1\leqslant\frac{n+\mathsf{def}_{r}(D)}{r}+2r-1,$
as desired. ∎
@@ 3. The Set-up
For the remainder of the paper, $G$ is a simple edge-colored graph with $n$ vertices, $n$ colors, and each color class of size at least $2$ and at most $r$, for $r\geq 2$. In other words, $G$ satisfies the hypotheses of Theorem 1.7. Let $k:=\lceil 78r\log{r}\rceil$ and $f(r)=80k^{2}r^{2}$. We will show that, for fixed $r$, we may take $\alpha_{r}:=(4r+2)f(r)+2r$ in Theorem 1.7.
To prove Theorem 1.7, we proceed by induction on $n$. The base case is $n\leq\alpha_{r}$, which is true since there is always a rainbow cycle of length at most $n$. Thus, for the remainder of the paper, we also assume that $n>\alpha_{r}$.
@@
Let $R=X\cup\{u_{1},\cdots,u_{m}\}\subseteq V(G)$, $M:=\{M_{u_{i}}:i\in[m]\}$ be a galaxy rooted at $\{u_{1},\cdots,u_{m}\}$, $C_{X}\subseteq C(X)$, and $d\in\mathbb{R}$ be such that
@@
    for every pair of vertices $u,v\in X$, there exists a rainbow path consisting of colors in $C_{X}$ from $u$ to $v$ in $G[X]$ of length at most $d$.
@@
$\frac{m+\sum_{i=1}^{m}(r-\lvert u_{i}\rvert_{M})}{r}+d+2.$
@@
See Figure 1 for an example with $m=7$. For each $u\in R$, we define a function $p:R\to R$ as follows. If $u\in X$, then $p(u)=u$. If $u\notin X$, but some $M$-neighbor $v$ of $u$ is in $X$, we arbitrarily choose one such $v$, and define $p(u)=v$. Otherwise, define $p(u)=u_{i}$, where $i$ is the minimum index such that $u_{i}$ is an $M$-neighbor of $u$. For example, in Figure 1, we have $p(u_{4})=u_{2}$.
@@
For the remaining case, let $u^{\prime}\in X$ and $v^{\prime}\in X$ be the other ends of $P^{u}$ and $P^{v}$. Let $Q^{u}=P^{u}\setminus u^{\prime}$ and $Q^{v}=P^{v}\setminus v^{\prime}$, and let $S^{u}$ be $Q^{u}$ minus its last vertex and $S^{v}$ be $Q^{v}$ minus its last vertex. If both $u$ and $v$ are in $X$, then we are done by (3). Thus, by symmetry, we may assume that $u=u_{a}$ for some $a\in[m]$ and that either $v\in X$ or $v=u_{b}$ for some $b<a$. By the previous case, we may assume that for all $x\in V(Q^{u})$ and $y\in V(Q^{v})$, $x$ and $y$ do not have any common $M$-neighbors in $\{u_{1},\dots,u_{m}\}$. Therefore,
@@
By (3), there is a rainbow path $P$ consisting of colors in $C_{X}$ from $u^{\prime}$ to $v^{\prime}$ in $G[X]$ of length at most $d$. Since $C_{X}\cap C(M)=\emptyset$, $P^{u}\cup P^{v}\cup P$ is a rainbow path consisting of colors in $C_{X}\cup C(M)$ from $u$ to $v$ in $G[R]$ of length at most
$\frac{m+\sum_{i=1}^{m}(r-|u_{i}|_{M})}{r}+d+2,$
@@ 4. Many Non-Stars
Recall that $k=\lceil 78r\log{r}\rceil$, $f(r)=80k^{2}r^{2}$, and $\alpha_{r}=(4r+2)f(r)+2r$. In this section we consider the case where $|N|>f(r)$. We say that a subset $N^{\prime}\subseteq N$ covers a color class $A$ if every edge in $A$ is incident to at least one vertex in $N^{\prime}$.
@@
subsets of $N$ of size $k$ cover some color class. Therefore, there exists a subset $N^{\prime}\subseteq N$ of size $k$ such that $N^{\prime}$ does not cover any color class. By choosing an edge of each color with neither end in $N^{\prime}$ we obtain a rainbow subgraph $H$ of $G$ with $|E(H)|=n$ and $|V(H)|\leqslant|V(G)\setminus N^{\prime}|=n-k$. In particular, $H$ is excess-$k$. Since $r\geq 2$ and $n\geq k$ (otherwise we have a rainbow cycle of length at most $k<\alpha_{r}$),  Corollary 2.2 gives that $H$ has a rainbow cycle of length at most
@@
Now, let $u_{1},\cdots,u_{m}$ be a maximal sequence of vertices in $G\setminus X$ such that there is a star color class $A_{i}$ centred at $u_{i}$ with $V(A_{i})\setminus\{u_{i}\}\subseteq X\cup\{u_{1},\dots,u_{i-1}\}$. Note that $u_{1},\dots,u_{m}$ exist, since we allow $m=0$. Let $T^{\prime}=X\cup\{u_{1},\cdots,u_{m}\}$, $c_{i}$ denote the color of $A_{i}$, and $C^{\prime}(T^{\prime})=C(X)\cup\{c_{1},\cdots,c_{m}\}$. Let $d:=\frac{|X|+\sum_{c\in C(X)}(r-|c_{G}|)}{r}-5r+2$. By Lemma 3.1, for all $u,v\in T^{\prime}$ there exists a rainbow path in $G[T^{\prime}]$ from $u$ to $v$ consisting of colors in $C(X)\cup\{c_{1},\dots,c_{m}\}$ of length at most
$\frac{m+\sum_{i=1}^{m}(r-\lvert u_{i}\rvert_{M})}{r}+d+2$ $=\frac{m+\sum_{i=1}^{m}(r-\lvert u_{i}\rvert_{M})}{r}+\left(\frac{|X|+\sum_{c\in C(X)}(r-|c_{G}|)}{r}-5r+2\right)+2$
@@
Note that when adding $Y$ to $T^{\prime}$ the ‘rainbow diameter’ increases by at most $2$, regardless of the size of $Y$. Therefore, for all $u,v\in T^{\prime}\cup Y$, there exists a rainbow path from $u$ to $v$ in $G[T^{\prime}\cup Y]$ of length at most
$\left(\frac{|T^{\prime}|+\sum_{c\in C(T^{\prime})}(r-|c_{G}|)}{r}-5r+4\right)+2\leqslant\frac{|T^{\prime}\cup Y|+\sum_{c\in C(T^{\prime}\cup Y)}(r-|c_{G}|)}{r}-5r+6-\frac{|Y|}{r}.$
It follows that $|Y|\leqslant 4r-1$; otherwise $T^{\prime}\cup Y$ contradicts the maximality of $X$. Now, we contract $T^{\prime}$ to a single vertex, delete all edges with colors in $T^{\prime}$, and remove parallel edges of the same color. Let the resulting edge-colored graph be $G^{\prime}$. Note that by the construction of $T^{\prime}$, $G^{\prime}$ contains $|V(G^{\prime})|$ color classes of size at least $2$, and at most $14r-$ 1 color classes have less edges in $G^{\prime}$ than they do in $G$ (since $|\mathcal{A}|\leqslant 4r-1$). By induction on $n$, $G^{\prime}$ has a rainbow cycle $D$ of length at most
@@ 5. Few Non-Stars
Let $N=\{v_{1},v_{2},\cdots,v_{t}\}$ be an arbitrary fixed ordering of the vertices of $N$. Let $U=V(G)\setminus N$ and choose a galaxy $M:=(M_{u}:u\in U)$ rooted at $U$ by arbitrarily choosing a star class $M_{u}$ of color $c_{u}$ centred at $u$ for each $u\in U$. An edge $e\in E(G)$ is an $M$-edge if $e\in E(M_{u})$ for some $u\in U$. For $Y\subseteq V(G)$ we define $\gamma(Y)$ to be the set of star vertices $u\in V(G)\setminus Y$ such that at least one $M$-neighbor of $u$ is in $Y$.
@@
    $\sum_{i=1}^{j}d_{i}+|\gamma(\bigcup_{i=1}^{j}T_{i})|\leqslant(4r+1)j$.
@@
Clearly, $T_{j}$ satisfies Item 1 and Item 2 (by the maximality of $u_{1},\dots,u_{m}$). Let $C^{\prime}:=\{c_{u_{i}}:i\in[m]\}$. By Lemma 3.1, for any pair of vertices $u,v\in T_{j}$, there is a rainbow path of $M$-edges from $u$ to $v$ in $G[T_{j}]$ of length at most
$\frac{m+\sum_{c\in C^{\prime}}(r-|c_{G}|)}{r}+d+2$ $=\frac{m+\sum_{c\in C^{\prime}}(r-|c_{G}|)}{r}+\left(\frac{|X|+\sum_{c\in C_{M}(X)}(r-|c_{G}|)}{r}+|s(X)|\right)+2$
@@
$\sum_{i=1}^{j}d_{i}+|\gamma(\bigcup_{i=1}^{j}T_{i})|\leqslant(4r+1)j$.
@@
Note that if we add $Y$ to $T_{j}$, then the ‘rainbow diameter’ increases by at most $2$, regardless of the size of $Y$. Thus, $|Y|\leqslant 4r-1$; otherwise, $Y\cup T_{j}$ contradicts the maximality of $X$ (see Lemma 4.2 for an analogous argument with more detail). Moreover, by definition, $s(T_{j})\subseteq\gamma(\bigcup_{i=1}^{j-1}T_{i})$ and $\gamma(\bigcup_{i=1}^{j}T_{i})\subseteq Y\cup(\gamma(\bigcup_{i=1}^{j-1}T_{i})\setminus s(T_{j}))$. Thus, $|\gamma(\bigcup_{i=1}^{j}T_{i})|\leqslant|Y|+|\gamma(\bigcup_{i=1}^{j-1}T_{i})|-|s(T_{j})|$. It follows that,
$\sum_{i=1}^{j}d_{i}+|\gamma(\bigcup_{i=1}^{j}T_{i})|$ $\leqslant d_{j}+|Y|-|s(T_{j})|+\sum_{i=1}^{j-1}d_{i}+|\gamma(\bigcup_{i=1}^{j-1}T_{i})|$
$=2+|Y|+\sum_{i=1}^{j-1}d_{i}+|\gamma(\bigcup_{i=1}^{j-1}T_{i})|$
$\leqslant(4r+1)+\sum_{i=1}^{j-1}d_{i}+|\gamma(\bigcup_{i=1}^{j-1}T_{i})|$
@@
The remaining case is $V(G)\neq\bigcup_{i=1}^{t}T_{i}$. Let $W=V(G)\setminus\left(\bigcup_{i=1}^{t}T_{i}\right)$. Let $D$ be the digraph with vertex set $W$ and for each $v\in W$ and each $M$-neighbor $u$ of $v$ in $W$, put an arc from $v$ to $u$ in $D$. By Item 2, every vertex $v\in W$ has at least one of its $M$-neighbors in $W$. Therefore, $D$ has no sink. By Item 4, at most $(4r+1)f(r)$ vertices in $W$ have an $M$-neighbor outside of $W$. Therefore, $\mathsf{def}_{r}(D)\leqslant(r-1)(4r+1)f(r)$. We now apply Corollary 2.4 in $D$ to obtain a directed cycle $C$ in $D$ of length at most
$\frac{|W|+(r-1)(4r+1)f(r)}{r}+2r-1$ $\leqslant\frac{n}{r}+\alpha_{r}.$